# JEDI-linear: Fast and Efficient Graph Neural Networks for Jet Tagging on FPGAs

### v1

@@ Abstract
Graph Neural Networks (GNNs), particularly Interaction Networks (INs), have shown exceptional performance for jet tagging at the CERN High-Luminosity Large Hadron Collider (HL-LHC). However, their computational complexity and irregular memory access patterns pose significant challenges for deployment on FPGAs in hardware trigger systems, where strict latency and resource constraints apply. In this work, we propose JEDI-linear, a novel GNN architecture with linear computational complexity that eliminates explicit pairwise interactions by leveraging shared transformations and global aggregation. To further enhance hardware efficiency, we introduce fine-grained quantization-aware training with per-parameter bitwidth optimization and employ multiplier-free multiply-accumulate operations via distributed arithmetic. Evaluation results show that our FPGA-based JEDI-linear achieves 3.7 to 11.5 times lower latency, up to 150 times lower initiation interval, and up to 6.2 times lower LUT usage compared to state-of-the-art designs while also delivering higher model accuracy and eliminating the need for DSP blocks entirely. In contrast, state-of-the-art solutions consume over 8,700 DSPs. This is the first interaction-based GNN to achieve less than 60 ns latency and currently meets the requirements for use in the HL-LHC CMS Level-1 trigger system. This work advances the next-generation trigger systems by enabling accurate, scalable, and resource-efficient GNN inference in real-time environments. Our open-sourced templates will further support reproducibility and broader adoption across scientific applications.
@@ I Introduction
To maximize hardware efficiency on FPGAs, we apply fine-grained quantization-aware training with per-parameter bitwidth optimization. In contrast to uniform quantization schemes [7, 1], our approach allows each parameter to adopt a bitwidth tailored to its impact on model performance. This is particularly well-suited for fully unrolled designs with strict latency requirements, in which each operation is mapped to dedicated hardware unit and reuse of these units for arithmetic operations is avoided. As a result, we can assign custom bitwidths to each operation, enabling fine-grained trade-offs between accuracy and resource usage. This bit-level control fully exploits the flexibility of reconfigurable logic on FPGAs, reducing redundancy and improving both latency and hardware efficiency.
@@
Although this work focuses on jet tagging in high-energy physics, the underlying principles and optimization techniques, such as linearized graph processing, mixed-precision quantization, and fully unrolled hardware design, are broadly applicable. We believe JEDI-linear can be generalized to other real-time applications where latency, throughput, and hardware efficiency are critical.
@@ III-A Global Information Gathering
Here, $f_{R}:\mathbb{R}^{2P}\rightarrow\mathbb{R}^{D_{E}}$ is a learnable function, which is implemented as a neural network in the original work [3]. Since $f_{R}$ is applied to every pair of particles, the overall computational complexity of this step is $\mathcal{O}(N_{O}^{2}\cdot C_{f_{R}})$, where $C_{f_{R}}$ denotes the cost of evaluating $f_{R}$ once. This quadratic scaling becomes a significant bottleneck when $N_{O}$ is even moderately large. For instance, $N_{O}=64$ would lead to over 4,000 interactions, making it infeasible for real-time applications. As a result, the pairwise interaction step dominates the inference cost and poses significant challenges for low-latency applications and hardware-constrained deployment environments, such as real-time triggering on FPGA-based systems.
Figure 2: Conventional interaction information gathering in JEDI-net.
@@ III-B The Proposed Global Information Gathering
Several previous studies have attempted to address the computational bottleneck inherent in the JEDI-net formulation when deployed on FPGAs. In [13], a custom MMM module is proposed for GNN feature transformation. By leveraging the sparse structures and binary features within the adjacency matrices $R_{R}$ and $R_{S}$, it can pick out elements and suppress unnecessary computations instead of conducting real multiplication operations. In [14], an outer-product-based MMM approach is introduced using a coarse-grained pipeline to improve throughput. More recently, [4] proposes a fused Edge-Node unit that integrates edge and node computations into a single hardware module. However, all of these designs still involve explicit edge-level computations, resulting in poor scalability and significant hardware resource consumption.
@@
$\begin{split}\bar{E}_{i}&=\sum_{j}f_{R}(I_{i}\|I_{j})\\ &=\sum_{j\neq i}\left(W_{1}I_{i}+W_{2}I_{j}+C\right)\\ &=W_{2}\sum_{j}\left(I_{j}\right)-I_{i}+\left(N_{O}-1\right)\left(W_{1}I_{i}+C\right)\end{split}$ (3)
@@
$\begin{split}\bar{E}_{i}^{\prime}&=\frac{1}{N_{O}}\bar{E}_{i}\\ &=W_{2}\frac{1}{N_{O}}\sum_{j}\left(I_{j}\right)-\frac{I_{i}}{N_{O}}+\frac{N_{O}-1}{N_{O}}\left(W_{1}I_{i}+C\right)\\ &\approx W_{2}\frac{1}{N_{O}}\sum_{j}\left(I_{j}\right)+W_{1}I_{i}+C\end{split}$ (4)
@@
Figure 3: The proposed interaction information gathering for JEDI-net.
@@ V-C Performance Analysis and Discussion
Compared to existing designs, our JEDI-linear models demonstrate a significant improvement in latency and initiation interval across all configurations. For example, with 16 input features and 64 particles, our design achieves an accuracy of 81.4% at only 79 ns latency, outperforming models such as GNN U4 and U5 [4] and GNN (FPL’22) [14] which require latencies of 905 ns and 10,660 ns, respectively, for higher accuracies. Fig. 8 shows that JEDI-linear achieves lower latency across various jet sizes compared to prior designs. Even at the largest configuration with 128 particles, our latency remains as low as 110 ns, which is suitable for hardware trigger systems that demand ultra-fast inference. In addition, our designs achieve a one-cycle initiation interval, a feat never before achieved by previous GNN designs.
Figure 8: Latency vs. number of particles per jet. Our JEDI-linear designs maintain low latency across input settings.
@@ VI Conclusions and Future Work
This work introduces JEDI-linear, a novel GNN architecture that enables linear-complexity and low-latency jet tagging on FPGAs. By eliminating explicit pairwise interactions and applying fine-grained quantization and distributed arithmetic, our designs achieve up to 82.4% classification accuracy, reduce inference latency to below 60 ns, and eliminate DSP usage entirely. Our work demonstrates that with careful algorithm-hardware co-design, high-performance GNNs can be deployed in real-time systems. In the future, we plan to explore broader GNN variants and automated co-design frameworks and extend JEDI-linear to other scientific domains.

### v2

@@ Abstract
Graph Neural Networks (GNNs), particularly Interaction Networks (INs), have shown exceptional performance for jet tagging at the CERN High-Luminosity Large Hadron Collider (HL-LHC). However, their computational complexity and irregular memory access patterns pose significant challenges for deployment on FPGAs in hardware trigger systems, where strict latency and resource constraints apply. In this work, we propose JEDI-linear, a novel GNN architecture with linear computational complexity that eliminates explicit pairwise interactions by leveraging shared transformations and global aggregation. To further enhance hardware efficiency, we introduce fine-grained quantization-aware training with per-parameter bitwidth optimization and employ multiplier-free multiply-accumulate operations via distributed arithmetic. Evaluation results show that our FPGA-based JEDI-linear achieves 3.7 to 11.5 times lower latency, up to 150 times lower initiation interval, and up to 6.2 times lower LUT usage compared to state-of-the-art GNN designs while also delivering higher model accuracy and eliminating the need for DSP blocks entirely. This is the first interaction-based GNN to achieve less than 60 ns latency and currently meets the requirements for use in the HL-LHC CMS Level-1 trigger system. This work advances the next-generation trigger systems by enabling accurate, scalable, and resource-efficient GNN inference in real-time environments. Our open-sourced templates will further support reproducibility and broader adoption across scientific applications.
@@ I Introduction
To maximize hardware efficiency on FPGAs, we apply fine-grained quantization-aware training with per-parameter bitwidth optimization. In contrast to uniform quantization schemes [7, 1], our approach allows each parameter to adopt a bitwidth tailored to its impact on model performance. This is particularly well-suited for fully unrolled designs with strict latency requirements, in which each operation is mapped to dedicated hardware unit and reuse of these units for arithmetic operations is avoided. As a result, we can assign custom bitwidths to each operation, enabling fine-grained trade-offs between accuracy and resource usage. This bit-level control fully exploits the flexibility of reconfigurable logic on FPGAs, improving both latency and hardware efficiency.
@@
Although this work focuses on jet tagging in high-energy physics, the underlying principles and optimization techniques, such as linearized processing, mixed-precision quantization, distributed arithmetic, and fully unrolled hardware design, are broadly applicable. They can be adapted for trustworthy DNNs [9] as well as low-latency variational autoencoders (VAEs) [10], transformers [11], and LLMs [12].
@@ III-A Global Information Gathering
Here, $f_{R}:\mathbb{R}^{2P}\rightarrow\mathbb{R}^{D_{E}}$ is a learnable function, which is implemented as a neural network in the original work [3]. $D_{E}$ denotes the edge latent dimension. Since $f_{R}$ is applied to every pair of particles, the overall computational complexity of this step is $\mathcal{O}(N_{O}^{2}\cdot C_{f_{R}})$, where $C_{f_{R}}$ denotes the cost of evaluating $f_{R}$ once. This quadratic scaling becomes a significant bottleneck when $N_{O}$ is even moderately large. For instance, $N_{O}=64$ would lead to over 4,000 interactions, making it infeasible for real-time applications. As a result, the pairwise interaction step dominates the inference cost and poses significant challenges for low-latency applications and hardware-constrained deployment environments, such as real-time triggering on FPGA-based systems.
Figure 2: Conventional interaction information gathering in JEDI-net. $R_{R}$ and $R_{S}$ are receiving and sending matrices; $N_{O}$ and $N_{E}$ are the numbers of particles and edges. $B1,B2,B,E,E,\bar{E}$ are all intermediate variables.
@@ III-B The Proposed Global Information Gathering
Several previous studies have attempted to address the computational bottleneck inherent in the JEDI-net formulation when deployed on FPGAs. In [16], a custom MMM (matrix-matrix multiplication) module is proposed for GNN feature transformation. By leveraging the sparse structures and binary features within the adjacency matrices $R_{R}$ and $R_{S}$, it can pick out elements and suppress unnecessary computations instead of conducting real multiplication operations. In [17], an outer-product-based MMM approach is introduced using a coarse-grained pipeline to improve throughput. More recently, [4] proposes a fused Edge-Node unit that integrates edge and node computations into a single hardware module. However, all of these designs still involve explicit edge-level computations, as shown in Fig. 2, resulting in poor scalability and significant hardware resource consumption.
@@
$\begin{split}\bar{E}_{i}&=\sum_{j\neq i}f_{R}(I_{i}\|I_{j})\\ &=\sum_{j\neq i}\left(W_{1}I_{i}+W_{2}I_{j}+C\right)\\ &=W_{2}\sum_{j}I_{j}-W_{2}I_{i}+\left(N_{O}-1\right)\left(W_{1}I_{i}+C\right)\end{split}$ (3)
@@
$\begin{split}\bar{E}_{i}^{\prime}&=\frac{1}{N_{O}}\bar{E}_{i}\\ &=W_{2}\frac{1}{N_{O}}\sum_{j}I_{j}-\frac{W_{2}I_{i}}{N_{O}}+\frac{N_{O}-1}{N_{O}}\left(W_{1}I_{i}+C\right)\\ &\approx W_{2}\frac{1}{N_{O}}\sum_{j}I_{j}+W_{1}I_{i}+C\end{split}$ (4)
@@
Figure 3: The proposed interaction information gathering for JEDI-net. Einsum denotes Einstein summation.
@@ V-C Performance Analysis and Discussion
Compared to existing designs, our JEDI-linear models demonstrate a significant improvement in latency and initiation interval across all configurations. For example, with 16 input features and 64 particles, our design achieves an accuracy of 81.4% at only 79 ns latency, outperforming models such as GNN U4 and U5 [4] and GNN (FPL’22) [17] which require latencies of 905 ns and 10,660 ns, respectively, for higher accuracies. Table I also shows that JEDI-linear achieves lower latency across various jet sizes compared to prior designs. Even at the largest configuration with 128 particles, our latency remains as low as 110 ns, which is suitable for hardware trigger systems that demand ultra-fast inference. In addition, our designs achieve a one-cycle initiation interval, a feat never before achieved by previous GNN designs.
@@ VI Conclusions and Future Work
This work introduces JEDI-linear, a novel GNN architecture that enables linear-complexity and low-latency jet tagging on FPGAs. By eliminating explicit pairwise interactions and applying fine-grained quantization and distributed arithmetic, our designs achieve up to 82.4% classification accuracy, reduce inference latency to below 60 ns, and eliminate DSP usage entirely. Our work demonstrates that with careful algorithm-hardware co-design, high-performance GNNs can be deployed in real-time systems. In the future, we plan to explore broader GNN variants, automation using metaprogramming (e.g., MetaML [32, 33]) and extend JEDI-linear to other scientific domains.